Akua is a school of Agriculture graduate and consumes 2 products:yoghurt and bread.Akua also earns a typical student income from the parent .400 cedis a month.She can either spend all on yoghurt and get 50,or she can spend it all on bread and get 200 loaves. Given the information construct the equation for Akua's budget line


1
Expert's answer
2021-05-05T14:34:49-0400

Solution:

Budget line: M = PxX + PyY

Where: M = Income

            X = Yoghurt

            Y = Bread

            Px = Price of yoghurt

            Py = Price of bread

Income = 400

Derive the price per unit for both yoghurt and bread.


Price per unit for yoghurt = "\frac{400}{50}" = 8


Price per unit for bread = "\frac{400}{200}" = 2

Px = 8

Py = 2

Therefore, the budget line is as follows:

M = PxX + PyY

400 = 8X + 2Y